# Bayesian Model Averaging in Causal Instrumental Variable Models

### v4

@@ 8 Conclusion
Whenever we use purely data-driven instrument selection in our simulations, we also allow naive BMA and OLS to include all instruments in the outcome model. Letting gIVBMA use the instrument information in the outcome model while the naive methods cannot would unfairly skew the comparison towards gIVBMA. This is particularly relevant when some of the instruments may be invalid.
S.4.3 Many weak instruments experiment
We now consider a setting with many individually weak instruments and $l=1$. Our simulation setup is similar to model C in Kuersteiner and Okui, (2010), where half of the instruments are relevant but increasingly weaker, and the other half is irrelevant. The main difference from their design is that we also consider exogenous covariates. We consider $p=20$ instruments and $k=10$ covariates generated from independent standard normal distributions and two different sample sizes $n\in\{50,500\}$. Then, we multiply all instruments and covariates with even indices by $100$ to have more variation in the variables’ scale. We expect IVBMA to struggle as its prior variance does not account for the scale, while the $g$-prior used in gIVBMA is scale-invariant. The instruments’ coefficients are generated by $\delta_{i}=c(p)\left(1-\frac{i}{p/2+1}\right)^{4}$ for $i=1,\dots,p/2$, and all other coefficients are set to zero. The constant $c(p)$ is chosen so that the first-stage $R^{2}$ of the instruments is approximately $0.01$ or $0.1$. The coefficients of all instruments with even indices are then divided by $100$ to adjust for the scaling. The covariates’ coefficients are set to $1/10$ or $1/1000$ for the first $k/2$ odd and even covariates, respectively, and zero else, and the treatment effect is set to $\tau=1/10$. The covariance matrix $\bm{\Sigma}$ is chosen to have $\sigma_{yy}=\sigma_{xx}=1$ and $\sigma_{yx}=1/2$.
We compare gIVBMA’s performance to naive BMA (not accounting for endogeneity), IVBMA, OLS, TSLS on the full model, oracle TSLS (O-TSLS) where we know the correct model specifications, which is of course not attainable in empirical work, TSLS with a jackknife-type fitted value in the first stage (JIVE, Angrist et al.,, 1999), TSLS with a jackknife and ridge-regularised first-stage (RJIVE, Hansen and Kozbur,, 2014), the post-LASSO estimator by Belloni et al., (2012) and the model-averaged TSLS estimator (MATSLS) with unrestricted weights by Kuersteiner and Okui, (2010). For gIVBMA, we consider three different $g$-prior specifications: BRIC, hyper-$g/n$, and the two-component prior (2C) with a hyper-$g/n$ prior on the covariate component $g_{C}$ and $c=1/2$. We do not use data-driven instrument selection in this example, but do not allow the instruments to enter the outcome model. Otherwise, we cannot use the two-component prior.
$n=50$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.5 0.5 0.06 1.36 0.44 0.44 0.08 1.38
gIVBMA (BRIC) 0.24 0.23 1.0 1.35 0.24 0.21 1.0 1.37
gIVBMA (hyper-$g/n$) 0.16 0.07 1.0 1.36 0.14 0.06 0.99 1.38
gIVBMA (2C) 0.2 0.08 1.0 1.36 0.15 0.05 0.99 1.39
IVBMA (KL) 0.54 0.54 0.83 9.71 0.34 0.34 0.91 6.26
OLS 0.52 0.52 0.05 1.43 0.45 0.45 0.07 1.47
TSLS 0.5 0.5 0.3 1.43 0.39 0.39 0.37 1.48
O-TSLS 0.46 0.46 0.63 1.39 0.36 0.36 0.63 1.41
JIVE 0.78 0.68 0.86 1.91 0.73 0.56 0.92 1.93
RJIVE 0.58 0.51 0.79 1.8 0.67 0.51 0.86 1.9
MATSLS 0.89 0.34 0.95 2.18 0.41 0.18 0.87 1.77
Post-LASSO 0.58 0.51 1.0 - 0.61 0.35 1.0 -
$n=500$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.5 0.5 0.0 1.28 0.45 0.45 0.0 1.29
gIVBMA (BRIC) 0.26 0.15 0.95 1.28 0.12 0.05 0.99 1.28
gIVBMA (hyper-$g/n$) 0.32 0.3 0.88 1.28 0.13 0.06 0.94 1.28
gIVBMA (2C) 0.41 0.33 0.81 1.28 0.12 0.04 0.93 1.28
IVBMA (KL) 0.26 0.26 0.67 76.94 0.09 -0.07 0.76 188.29
OLS 0.5 0.5 0.0 1.28 0.45 0.45 0.0 1.3
TSLS 0.39 0.39 0.45 1.31 0.15 0.14 0.75 1.37
O-TSLS 0.33 0.33 0.66 1.33 0.08 0.06 0.92 1.4
JIVE 0.96 0.79 0.82 1.72 0.18 -0.15 0.98 1.57
RJIVE 0.97 0.74 0.82 1.75 0.18 -0.16 0.98 1.57
MATSLS 0.34 0.02 0.87 1.62 0.1 0.01 0.96 1.45
Post-LASSO 1.13 0.26 1.0 - 0.11 0.02 0.99 -
Table S.1: Many Weak Instruments: MAE, bias, coverage, and mean LPS on 100 simulated datasets with many weak instruments. The best values in each column are printed in bold. Post-Lasso only returns estimates for $\tau$ but not for the other coefficients, so we cannot compute the LPS. When no instrument is selected, no effect estimates are provided, therefore, we do not consider those cases. The number of times no instruments were selected by Post-Lasso in the first stage is (from top-left to bottom-right): 93, 85, 74, 1.
Table S.1 shows the results. The gIVBMA methods do best overall in terms of LPS and in the small sample, also in terms of MAE and median bias. Their coverage is too high for small $n$, while the hyper-$g/n$ specifications slightly undercover for larger $n$. In some cases, TSLS and O-TSLS come close, but these can also perform very badly, in particular with low coverage. OLS and naive BMA are not too far from TSLS in terms of MAE and bias, but do considerably worse with $n=500$ and stronger instruments. In line with the discussion before the start of this subsection, the predictive performance of BMA (and OLS in the larger sample) is comparable to gIVBMA, but coverage is far too low. The jackknife-based methods tend to have higher MAE and bias but better coverage than regular TSLS. MATSLS performs very well overall in the $n=500$ case but also has good coverage for $n=50$. Post-Lasso performs well with the stronger instruments. However, only in the $n=500$ scenario with stronger instruments does the Lasso tend to select instruments, whereas in the other cases, it often does not select any. IVBMA undercovers in all four scenarios, and in the smaller sample, it also performs substantially worse than gIVBMA in terms of MAE and bias. IVBMA’s predictive performance is very poor in all scenarios (especially for $n=500$) since the prior on the regression coefficients is not scale-invariant.
$n=50$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.06 0.06 0.32 1.47 0.05 0.05 0.39 1.44
gIVBMA (BRIC) 0.02 0.01 0.95 1.5 0.02 -0.0 0.99 1.51
gIVBMA (hyper-$g/n$) 0.02 0.0 0.94 1.51 0.02 -0.01 0.98 1.51
gIVBMA (2C) 0.02 0.0 0.95 1.51 0.02 -0.01 0.97 1.51
IVBMA (KL) 0.11 0.06 0.53 14.84 0.1 -0.04 0.6 12.64
OLS 0.06 0.06 0.28 1.6 0.05 0.05 0.36 1.51
TSLS 0.07 0.07 0.61 1.62 0.05 0.05 0.6 1.51
O-TSLS 0.07 0.06 0.77 1.59 0.04 0.04 0.87 1.47
JIVE 0.11 0.05 0.98 2.07 0.14 0.05 1.0 2.13
RJIVE 0.14 0.05 1.0 2.69 0.12 0.07 1.0 2.62
MATSLS 0.17 0.05 0.98 2.49 0.08 0.03 0.92 1.91
Post-LASSO 0.08 0.08 1.0 - 0.15 0.01 0.75 -
$n=500$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.06 0.06 0.0 1.35 0.05 0.05 0.0 1.37
gIVBMA (BRIC) 0.01 -0.0 0.94 1.42 0.01 -0.0 0.95 1.44
gIVBMA (hyper-$g/n$) 0.01 -0.0 0.94 1.42 0.01 -0.0 0.95 1.44
gIVBMA (2C) 0.01 -0.0 0.95 1.42 0.01 -0.0 0.94 1.44
IVBMA (KL) 0.1 -0.05 0.29 216.37 0.08 -0.05 0.54 140.27
OLS 0.06 0.06 0.0 1.36 0.05 0.05 0.0 1.38
TSLS 0.04 0.04 0.69 1.38 0.02 0.02 0.81 1.41
O-TSLS 0.04 0.03 0.85 1.4 0.02 0.01 0.9 1.42
JIVE 0.1 0.08 0.83 1.64 0.05 -0.03 0.97 1.79
RJIVE 0.09 0.07 0.89 1.69 0.09 0.0 0.97 1.99
MATSLS 0.08 0.01 0.98 1.75 0.02 -0.0 0.94 1.48
Post-LASSO 0.42 -0.28 1.0 - 0.02 0.0 1.0 -
Table S.2: Many Weak Instruments: MAE, bias, coverage, and mean LPS with a Poisson endogenous variable and many weak instruments based on 100 simulated datasets. The best values in each column are printed in bold. Post-Lasso only returns estimates for $\tau$, but not for the other coefficients, so we cannot compute the LPS. When no instrument is selected, no effect estimates are provided, therefore, we do not consider those cases. The number of times no instruments were selected in the first stage is (from top-left to bottom-right): 98, 96, 95, 48.
Table S.2 presents the results, which are qualitatively somewhat similar to the ones above. Notably, the naive methods (OLS and regular BMA) are much closer than in the Gaussian setting. This is perhaps a consequence of the additional variance induced by the non-linear transformation, diluting the degree of endogeneity. Also, IVBMA’s small sample performance in terms of MAE and bias is much better, but its predictive performance is still very poor. Post-Lasso has even more instances of not selecting a single instrument, so its performance is based on very few iterations.
S.4.4 Multiple endogenous variables with correlated instruments
@@
Table S.4 presents the median posterior inclusion probabilities from the simulation with multiple endogenous variables. The gIVBMA variants always include the correct variables while only very sparsely including wrong instruments in the smaller sample and not at all in the larger sample. IVBMA correctly selects the included instruments, but also includes other unnecessary variables too often for the first endogenous variable. For the second endogenous variable, some of the true instruments are hardly included in the $n=50$ scenario, while the performance is good for $n=500$.
@@
BMA (hyper-$g/n$) 0.48 0.42 0.47 0.92 1.39
gIVBMA (BRIC) 0.24 0.06 0.93 1.0 1.23
gIVBMA (hyper-$g/n$) 0.22 0.04 0.92 0.99 1.23
IVBMA (KL) 0.52 0.5 0.92 0.94 1.26
OLS 0.95 0.65 0.0 0.93 1.46
TSLS 0.28 0.16 0.86 0.98 1.46
O-TSLS 0.27 0.09 0.89 0.98 1.47
MATSLS 0.46 0.13 0.91 0.95 1.57
@@
BMA (hyper-$g/n$) 0.76 0.66 0.0 0.99 1.14
gIVBMA (BRIC) 0.08 0.01 0.99 0.87 1.15
gIVBMA (hyper-$g/n$) 0.1 0.01 0.94 0.89 1.15
IVBMA (KL) 0.25 0.23 0.98 0.92 1.14
OLS 0.77 0.67 0.0 0.97 1.16
TSLS 0.11 0.01 0.95 0.94 1.43
O-TSLS 0.11 0.02 0.96 0.95 1.43
MATSLS 0.1 0.01 0.94 0.94 1.43
Table S.3: Multiple endogenous variables with correlated instruments: MAE, bias, coverage, and mean LPS based on 100 simulated datasets with two endogenous variables (one Gaussian and one Beta). The best values in each column are in bold.
gIVBMA (BRIC) gIVBMA (hyper-$g/n$) IVBMA (X1) IVBMA (X2)
@@
S.5.1 Returns to schooling
Here, we present additional details for the returns to schooling example based on the Card, (1995) dataset. Table S.5 includes definitions of all the variables used. Tables S.6 and S.7 present the posterior inclusion probabilities obtained with the Card, (1995) dataset, including and excluding parental education, respectively, by gIVBMA, IVBMA, and naive BMA (only the outcome model). Figure S.2 and Table S.8 present the posterior results and inclusion probabilities for the model without parental education on the smaller shared set of observations. Figure S.3 shows the posterior distributions of the treatment residual variance on all three subsets of the dataset.
@@
gIVBMA (hyper-$g/n$) gIVBMA (BRIC) IVBMA BMA (hyper-$g/n$)
@@
age 0.742 0.176 0.492 0.533 0.989 1.0 0.89
agesq 0.262 0.174 0.536 0.534 0.013 1.0 0.352
nearc2 0.071 0.062 0.385 0.08 0.343 0.304 0.672
nearc4 0.001 0.298 0.032 0.572 0.069 0.893 0.091
momdad14 0.989 1.0 0.906 1.0 0.997 1.0 0.104
sinmom14 0.003 0.008 0.012 0.01 0.032 0.132 0.082
step14 0.001 0.006 0.009 0.011 0.059 0.148 0.124
black 0.035 1.0 0.167 1.0 0.223 1.0 1.0
south 0.005 1.0 0.177 0.995 0.125 0.998 1.0
smsa 0.006 1.0 0.152 1.0 0.063 1.0 1.0
married 1.0 1.0 1.0 1.0 1.0 1.0 1.0
reg662 0.001 0.004 0.015 0.015 0.033 0.128 0.089
reg663 0.206 0.052 0.632 0.054 0.669 0.309 0.757
reg664 0.006 0.005 0.022 0.016 0.09 0.141 0.21
reg665 0.0 0.005 0.016 0.016 0.03 0.079 0.147
reg666 0.0 0.004 0.016 0.01 0.031 0.109 0.089
reg667 0.0 0.004 0.009 0.018 0.024 0.098 0.085
reg668 0.395 0.027 0.665 0.126 0.938 0.713 0.922
reg669 0.004 0.038 0.047 0.111 0.06 0.536 0.184
Table S.6: Returns to schooling: Posterior inclusion probabilities for the Card, (1995) dataset excluding parental education, using the full sample of $n=3,003$ observations.
gIVBMA (hyper-$g/n$) gIVBMA (BRIC) IVBMA BMA (hyper-$g/n$)

### v5

@@ 8 Conclusion
Whenever we use purely data-driven instrument selection in our simulations, we also allow naive BMA to include all instruments in the outcome model. Letting gIVBMA use the instrument information in the outcome model while the naive method cannot would unfairly skew the comparison towards gIVBMA. This is particularly relevant when some of the instruments may be invalid.
S.5.3 Invalid Instruments
Figure S.3 displays traceplots of the included variables in both models and the number of implied instruments at each iteration of the chain for a single simulated dataset. We only consider the hyper-$g/n$ specifications with the full inverse Wishart prior and the Cholesky prior with $\omega_{a}=0.1$, as these perform best in the simulation experiment (see results in Section 6.1 of the main text). For the smaller sample size $n=50$, the chains still visit non-identified models with high frequency (in line with the averages over 100 datasets in Table 1 in the main text), but mix well. For $n=500$, the chains mix well and spend most of their time at $N_{Z}=6$ or the true value, $N_{Z}=7$. The difference in behaviour between the small and large sample is mainly driven by the treatment model. In the small sample, the treatment model includes too few instruments in most iterations. This is a consequence of the parsimony preference of Bayes factors and the small amount of data information.
Figure S.3: Invalid Instruments: A chain of the outcome model size ($L$), treatment model size ($M$), and the number of valid and relevant instruments $N_{Z}$ implied by outcome and treatment model at each iteration for both considered gIVBMA variants. These results are based on a single simulated dataset of size $n=50$ (left) and $n=500$ (right), respectively, with $s=3$ out of $10$ instruments being invalid.
S.5.4 Multiple endogenous variables with correlated instruments
@@
Posterior results on MAE, bias, coverage and prediction are summarized in Table S.1, which is discussed in the main text.
Table S.2 presents the median posterior inclusion probabilities from the simulation with multiple endogenous variables. The gIVBMA variants always include the correct variables while only very sparingly including wrong instruments in the smaller sample and not at all in the larger sample. IVBMA correctly selects the included instruments, but also includes other unnecessary variables too often for the Gaussian endogenous variable. For the second (Beta) endogenous variable, some of the true instruments are often excluded in the $n=50$ scenario, while the performance is good for $n=500$.
@@
BMA (hyper-$g/n$) 0.53 0.34 0.64 0.9 1.38
gIVBMA (hyper-$g/n$, IW) 0.27 0.05 0.97 0.97 1.19
gIVBMA (hyper-$g/n$, $\omega_{a}=0.1$) 0.25 0.05 0.97 0.97 1.18
IVBMA 0.42 0.39 0.98 1.0 1.29
TSLS 0.3 0.16 0.96 0.94 1.46
O-TSLS 0.32 0.12 0.98 0.96 1.48
MATSLS 0.49 0.18 0.95 0.95 1.65
@@
BMA (hyper-$g/n$) 0.76 0.67 0.0 0.95 1.13
gIVBMA (hyper-$g/n$, IW) 0.08 0.02 0.92 0.92 1.14
gIVBMA (hyper-$g/n$, $\omega_{a}=0.1$) 0.1 0.02 0.91 0.96 1.13
IVBMA 0.21 0.18 0.98 0.93 1.14
TSLS 0.11 0.06 0.9 0.96 1.42
O-TSLS 0.11 0.04 0.91 0.96 1.42
MATSLS 0.1 0.05 0.89 0.95 1.42
Table S.1: Multiple endogenous variables with correlated instruments: Results with two endogenous variables (one Gaussian and one Beta) based on 100 simulated datasets. The best values in each column are printed in bold.
gIVBMA (h-$g/n$, IW) gIVBMA (h-$g/n$, $\omega_{a}=0.1$) IVBMA (X1) IVBMA (X2)
@@
S.5.5 Many weak instruments
We now consider a setting with many individually weak instruments and $l=1$. Our simulation setup is similar to model C in Kuersteiner and Okui, (2010), where half of the instruments are relevant but increasingly weaker, and the other half is irrelevant. The main difference from their design is that we also consider exogenous covariates. We consider $p_{1}=20$ instruments and $p_{2}=10$ covariates generated from independent standard normal distributions and two different sample sizes $n\in\{50,500\}$. Then, we multiply all instruments and covariates with even indices by $100$ to have more variation in the variables’ scale. We expect IVBMA to struggle as its prior variance does not account for the scale, while the $g$-prior used in gIVBMA is scale-invariant. The instruments’ coefficients are generated by
$\delta_{i}=c(p_{1})\left(1-\frac{i}{p_{1}/2+1}\right)^{4},\quad i=1,\dots,p_{1}/2,$
and all other coefficients are set to zero. The constant $c(p_{1})$ is chosen so that the first-stage $R^{2}$ of the instruments is approximately $0.01$ or $0.1$. The coefficients of all instruments with even indices are then divided by $100$ to adjust for the scaling. The covariates’ coefficients are set to $1/10$ or $1/1000$ for the first $p_{2}/2$ odd and even covariates, respectively, and zero else, and the treatment effect is set to $\tau=1/10$. The covariance matrix $\bm{\Sigma}$ is chosen to have $\sigma_{yy}=\sigma_{xx}=1$ and $\sigma_{yx}=1/2$.
We compare gIVBMA’s performance to naive BMA (not accounting for endogeneity), IVBMA, Bayesian IV with a horseshoe prior, TSLS on the full model, oracle TSLS (O-TSLS) where we know the correct model specifications, which is of course not attainable in empirical work, TSLS with a jackknife-type fitted value in the first stage (JIVE, Angrist et al.,, 1999), TSLS with a jackknife and ridge-regularised first-stage (RJIVE, Hansen and Kozbur,, 2014), the post-LASSO estimator by Belloni et al., (2012), and the model-averaged TSLS estimator (MATSLS) with unrestricted weights by Kuersteiner and Okui, (2010). For gIVBMA, we consider three different $g$-prior specifications: hyper-$g/n$ with the full inverse Wishart prior and the Cholesky-based prior with $\omega_{a}=0.1$, and the two-component prior (2C) with a hyper-$g/n$ prior on the covariate component $g_{C}$ and $c=1/2$ and the full inverse Wishart covariance prior. In this setting, the instruments are assumed to be valid a priori and cannot be included in the outcome model, as described in Section S.2.
$n=50$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.46 0.46 0.08 1.38 0.45 0.45 0.08 1.48
gIVBMA (IW) 0.17 0.08 0.93 1.39 0.16 0.06 1.0 1.48
gIVBMA ($\omega_{a}=0.1$) 0.36 0.36 0.85 1.38 0.37 0.37 0.89 1.48
gIVBMA (IW, 2C) 0.19 0.07 0.92 1.38 0.2 0.07 1.0 1.48
BayesHS 0.03 0.03 0.93 2.2 0.03 0.03 1.0 2.23
TSLS 0.47 0.47 0.29 1.47 0.44 0.44 0.36 1.56
O-TSLS 0.51 0.51 0.48 1.43 0.37 0.37 0.7 1.5
JIVE 0.76 0.47 0.86 1.89 0.83 0.52 0.91 2.01
RJIVE 0.58 0.5 0.81 1.84 0.71 0.54 0.89 2.04
MATSLS 0.79 0.36 0.9 2.03 0.47 0.02 0.88 1.98
IVBMA 0.13 0.11 0.99 1.41 0.11 0.05 0.99 1.46
Post-LASSO 0.53 0.53 0.8 - 0.34 0.18 1.0 -
$n=500$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.49 0.49 0.0 1.28 0.45 0.45 0.0 1.3
gIVBMA (IW) 0.31 0.25 0.9 1.28 0.12 0.04 0.9 1.28
gIVBMA ($\omega_{a}=0.1$) 0.41 0.41 0.78 1.28 0.09 0.07 0.92 1.28
gIVBMA (IW, 2C) 0.34 0.32 0.84 1.28 0.12 0.02 0.93 1.28
BayesHS 0.02 0.02 1.0 2.14 0.02 0.02 1.0 2.14
TSLS 0.38 0.38 0.44 1.32 0.12 0.1 0.83 1.39
O-TSLS 0.34 0.33 0.65 1.35 0.09 0.04 0.9 1.41
JIVE 0.9 0.56 0.73 1.89 0.22 0.19 0.97 1.62
RJIVE 0.92 0.56 0.73 1.9 0.22 0.2 0.97 1.62
MATSLS 0.41 0.1 0.92 1.64 0.1 0.0 0.95 1.45
IVBMA 0.12 0.01 0.8 1.29 0.1 0.08 0.71 1.29
Post-LASSO 0.48 0.26 1.0 - 0.1 0.0 0.98 -
Table S.3: Many Weak Instruments: MAE, median bias, coverage, and mean LPS with a Gaussian endogenous variable based on 100 simulated datasets. The best values in each column are printed in bold. Post-Lasso only returns estimates for $\tau$, but not for the other coefficients, so we cannot compute the LPS. When no instrument is selected, no effect estimates are provided, therefore we do not consider those cases. The number of times no instruments were selected in the first stage is (from top-left to bottom-right): 95, 85, 76, 3.
Table S.3 shows the results. The full inverse Wishart with the regular hyper-$g/n$ prior is the best gIVBMA variant with good MAE and coverage in all scenarios. The performance of the two-component (2C) prior is very similar, while the Cholesky prior with $\omega_{a}=0.1$ is the worst variant of gIVBMA in three out of four scenarios. The Bayesian horseshoe does best in terms of MAE and bias, but overcovers in three out of four scenarios and predicts badly. IVBMA tends to do slightly better than gIVBMA in terms of MAE, but has bad coverage for $n=500$. TSLS and O-TSLS only come close in the larger sample size with stronger instruments. Naive BMA is not too far from TSLS in terms of MAE and bias, but does considerably worse with $n=500$ and stronger instruments. In line with the discussion at the beginning of Section 6, the predictive performance of BMA is comparable to gIVBMA, but coverage is far too low. The jackknife-based methods tend to have higher MAE and bias but better coverage than regular TSLS. MATSLS performs very well overall in the $n=500$ case but also has good coverage for $n=50$. Post-Lasso performs well with the stronger instruments. However, only in the $n=500$ scenario with stronger instruments does the Lasso tend to select instruments, whereas in the other cases, it often does not select any.
$n=50$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.064 0.064 0.32 1.42 0.049 0.049 0.39 1.47
gIVBMA (IW) 0.023 0.007 0.93 1.51 0.028 0.001 0.96 1.53
gIVBMA ($\omega_{a}=0.1$) 0.03 0.026 0.91 1.45 0.025 0.018 0.93 1.48
gIVBMA (IW, 2C) 0.026 0.003 0.94 1.5 0.028 0.002 0.94 1.52
BayesHS 0.021 0.02 0.97 2.2 0.018 0.017 0.99 2.17
TSLS 0.063 0.063 0.58 1.55 0.05 0.05 0.62 1.54
O-TSLS 0.063 0.061 0.75 1.55 0.045 0.044 0.79 1.54
JIVE 0.091 0.065 0.91 1.9 0.104 0.053 0.98 2.12
RJIVE 0.127 0.049 0.98 2.54 0.112 0.031 1.0 2.82
MATSLS 0.136 0.054 0.96 2.11 0.075 0.019 0.95 2.06
IVBMA 0.099 0.076 0.55 1.52 0.099 0.091 0.51 1.52
Post-LASSO 0.066 0.066 1.0 - 0.107 0.08 1.0 -
$n=500$
$R^{2}_{f}=0.01$ $R^{2}_{f}=0.1$
MAE Bias Cov. LPS MAE Bias Cov. LPS
BMA (hyper-$g/n$) 0.056 0.056 0.0 1.35 0.047 0.047 0.0 1.36
gIVBMA (IW) 0.007 0.001 0.94 1.42 0.006 0.0 0.96 1.42
gIVBMA ($\omega_{a}=0.1$) 0.007 0.003 0.93 1.41 0.006 0.003 0.94 1.42
gIVBMA (IW, 2C) 0.007 0.001 0.94 1.42 0.006 0.0 0.96 1.42
BayesHS 0.035 0.035 1.0 2.28 0.031 0.031 1.0 2.28
TSLS 0.052 0.052 0.61 1.38 0.02 0.019 0.84 1.4
O-TSLS 0.048 0.048 0.86 1.38 0.018 0.012 0.9 1.41
JIVE 0.079 0.062 0.88 1.71 0.071 0.018 1.0 1.78
RJIVE 0.066 0.053 0.89 1.7 0.097 0.018 0.98 1.86
MATSLS 0.089 0.013 0.98 1.85 0.021 0.003 0.95 1.46
IVBMA 0.1 0.066 0.33 1.36 0.082 0.041 0.46 1.38
Post-LASSO 0.358 0.121 1.0 - 0.019 0.013 1.0 -
Table S.4: Many Weak Instruments: MAE, median bias, coverage, and mean LPS with a Poisson endogenous variable based on 100 simulated datasets. The best values in each column are printed in bold. Post-Lasso only returns estimates for $\tau$, but not for the other coefficients, so we cannot compute the LPS. When no instrument is selected, no effect estimates are provided, therefore we do not consider those cases. The number of times no instruments were selected in the first stage is (from top-left to bottom-right): 97, 98, 95, 44.
Table S.4 presents the results, which are qualitatively somewhat similar to the ones above. Notably, the differences are smaller than in the Gaussian setting. This is perhaps a consequence of the additional variance induced by the non-linear transformation, diluting the degree of endogeneity. All gIVBMA variants display strong performance. Its MAE is only beaten by BayesHS in the small $n$ scenario, but its coverage and predictions are much better. IVBMA’s performance is now much worse, in particular its coverage is very low. Post-Lasso has even more instances of not selecting a single instrument, so its performance is based on very few iterations.
For a single realisation of the data-generating process with a Gaussian endogenous variable, Figure S.4 shows traceplots of the treatment and outcome model sizes. In the treatment model, the $n=50$ chains spend quite some time at the intercept only model, where no instruments are included. However, for $n=500$, the chains for the treatment models mix nicely with most of their mass between $5$ and $10$ included variables. The outcome models’ chains mix well between 5105-10 included covariates (the true model includes $5$) for $n=50$, whereas for the larger sample they concentrate on $5$ or $6$ included covariates.
Figure S.4: Many Weak Instruments: A chain of treatment (top) and outcome (bottom) model sizes of length $5,000$ (after $1,000$ burn in iterations) for the three considered gIVBMA variants. These results are based on a single simulated dataset for $n=50$ (left) and $n=500$ (right), respectively, with instrument strength of $R_{f}^{2}\approx 0.1$.
@@
S.6.1 Returns to schooling
Here, we present additional details for the returns to schooling example based on the Card, (1995) dataset. Table S.5 includes definitions of all the variables used.
Figure S.5 presents the posterior results on the dataset only including observations with complete parental education. Tables S.6 and S.7 present the posterior inclusion probabilities obtained with the full Card, (1995) dataset and the smaller complete dataset, respectively, by gIVBMA, IVBMA, and naive BMA (only the outcome model). Figure S.6 presents the implied posterior distributions of the number of valid and relevant instruments.
@@
fathmiss =1 if fatheduc is missing
mothmiss =1 if motheduc is missing
@@
gIVBMA (IW) gIVBMA ($\omega_{a}=0.1$) IVBMA BMA
@@
exper 1.0 1.0 1.0 1.0 0.006 1.0 1.0
expersq 1.0 0.0 1.0 0.014 0.0 0.003 1.0
nearc2 0.024 0.009 0.058 0.0 0.218 0.032 0.704
nearc4 0.002 0.971 0.0 0.984 0.048 1.0 0.092
momdad14 0.005 1.0 0.005 1.0 0.008 1.0 0.176
sinmom14 0.008 0.009 0.002 0.023 0.048 0.218 0.079
step14 0.0 0.003 0.014 0.01 0.333 0.208 0.164
black 1.0 1.0 1.0 1.0 1.0 1.0 1.0
south 1.0 0.041 1.0 0.073 1.0 0.192 1.0
smsa 1.0 0.927 1.0 0.9 1.0 0.991 1.0
married 1.0 0.982 1.0 0.958 1.0 0.989 1.0
reg662 0.0 0.009 0.002 0.0 0.074 0.143 0.113
reg663 0.101 0.0 0.093 0.007 0.086 0.104 0.782
reg664 0.048 0.0 0.013 0.023 0.581 0.477 0.205
reg665 0.002 0.014 0.019 0.003 0.032 0.164 0.174
reg666 0.0 0.03 0.0 0.0 0.033 0.17 0.148
reg667 0.0 0.002 0.0 0.004 0.023 0.173 0.108
reg668 0.771 0.087 0.773 0.064 0.893 0.687 0.947
reg669 0.0 0.31 0.0 0.295 0.056 0.853 0.22
fatheduc 0.0 1.0 0.007 1.0 0.0 1.0 0.14
motheduc 0.0 1.0 0.007 1.0 1.0 1.0 0.467
fathmiss 0.0 0.095 0.0 0.15 0.026 0.829 0.083
mothmiss 0.009 0.032 0.0 0.016 0.012 0.545 0.11
Table S.6: Returns to schooling: Posterior inclusion probabilities for the Card, (1995) dataset using the full (imputed) dataset ($n=3,003$).
Figure S.5: Returns to schooling: Posterior distributions of the treatment effect of education (Rao-Blackwellized for gIVBMA; Kernel density estimate for IVBMA) and the covariance ratio $\sigma_{yx}/\sigma_{xx}$ based on the Card, (1995) dataset only using the observations that have complete entries for parental education ($n=2,215$). The algorithm was run for $5,000$ iterations, the first $500$ of which were discarded as burn-in.
gIVBMA (IW) gIVBMA ($\omega_{a}=0.1$) IVBMA BMA